OpenStudy (anonymous):

Approximate the sum of the series (Sigma) 1/(n^3) by using the first 5 terms. Which is just: 1+1/8+1/27+1/64+1/125 = 1.18543.... How do I estimate the error involved in this approximation?

OpenStudy (amistre64):

the eror is never greater than the next term, or something along those lines

OpenStudy (anonymous):

So I would be comparing the error of 1.18543 and 1.18543 + 1/(6^3)?

OpenStudy (amistre64):

wish i could tell, i never really got that far with it :)

OpenStudy (amistre64):

http://ecademy.agnesscott.edu/~lriddle/apcalculus/approxSeries.pdf might be useful

OpenStudy (amistre64):

total sum = partial + remainder \[s=s_n+R_n\] \[R_n=s-s_n\] \[R_{max}=\int_{n+1}^{inf} f(x)dx\] \[R_{min}=\int_{n}^{inf} f(x)dx\]

OpenStudy (amistre64):

i think i might have those backwards tho

OpenStudy (amistre64):

anyhoos \[\int_{6}^{inf}x^{-3}dx=-\frac{1}{2(inf)^2}+\frac{1}{2(6)^2}\] \[\int_{5}^{inf}x^{-3}dx=-\frac{1}{2(inf)^2}+\frac{1}{2(5)^2}\] so it looks like the total sum rests between sn + 1/72 and sn+1/50

OpenStudy (amistre64):

well, sn being the 5th partial sum in this case